OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 3) and perpendicular to:

OpenStudy (anonymous):

\[y = -\frac{ 1 }{ 3 }x + 7\]

OpenStudy (anonymous):

Ok, first we need the slope of the unknown equation.

OpenStudy (anonymous):

Okay, I forget what to do.

OpenStudy (anonymous):

The perpendicular slope is the negative of the reciprocal

OpenStudy (anonymous):

So the unknown slope is \[\frac{3}{1}\]

OpenStudy (anonymous):

or just 3.

OpenStudy (anonymous):

yep, or just 3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

So so far the equation of this second line is y = 3x + b

OpenStudy (anonymous):

So now we just need b

OpenStudy (anonymous):

So, \[3 = 3(-1) + b\]\[3 = -3 + b\]\[6 = b\]?

OpenStudy (anonymous):

Oh yeah that works too

OpenStudy (anonymous):

I was just going to add the slope to the point

OpenStudy (anonymous):

Your way is better

OpenStudy (anonymous):

Hahah. XD So, my final answer is y = 3x + 6 right?

OpenStudy (anonymous):

Yep lol

OpenStudy (anonymous):

Haha, thank you broskii

OpenStudy (anonymous):

No prob

OpenStudy (anonymous):

Wanna help with another?

OpenStudy (anonymous):

Sure but open it as a new question

OpenStudy (anonymous):

Aight.